A marble rolls off of a table and travels a horizontal range of 0.35 m in the air before landing on the ground. If the table is 0.88 m tall, calculate the marble’s initial velocity as it leaves the table.


1
Expert's answer
2021-01-13T11:35:44-0500

Let's first find the time that the marble takes to reach the ground:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot 0.88\ m}{9.8\ \dfrac{m}{s^2}}}=0.42\ s."

Finally, we can find the initial velocity as it leaves the table:


"x=v_0t,""v_0=\dfrac{x}{t}=\dfrac{0.35\ m}{0.42\ s}=0.83\ \dfrac{m}{s}."

Answer:

"v_0=0.83\ \dfrac{m}{s}."